50.0 mL of 0.275 M HNO2 is titrated to its equivalence point with 1.00 M NaOH. What is the pH at the equivalence point?


Ka for HNO2 is 4.0x10^-4
Answer #1

pH formulas for salts weak acidt strong base ->salft to pH = Ź [pkw + pka + 10g c] strong acid & weakbase salf toto pH = ₃Kaz 4x104 Ang molarity 20 - 275 | pka, log ka = log ve x104) volume 50ml=0.052 1 pka_ 3.398 M = 2 no of moles 0.27500-05 l pkANO: molasity 20. 275 ka= 4x104 1 pka_ -log ka = log (24x154) volume som = 0.054 pka_ 3.398 M= A no of moles= 0.27580-05 / pk